What is the weight of a box that has a mass of 81.3 kg if it were placed of the Moon, whose gravitational attraction is 0.17 times that of the Earth?
Answer #1

Selmer weight on moon (lol) = mg moon Umoon Where Yoon = 0.17 Jeast = 0.17(9.80m/s) Jmoon ith = a 1.6677 m/s2 = mass=8103kg m